The Candy Emporium sells two types of candy- type A sells for $1.50 per pound and type B sells for $2.00 per pound. How kuch of each type of candy should be combined to obtain 20 pounds of a mixture which sells for $1.80 per pound? 
Answer by stanbon(75887)

The Candy Emporium sells two types of candy- type A sells for $1.50 per pound and type B sells for $2.00 per pound. How kuch of each type of candy should be combined to obtain 20 pounds of a mixture which sells for $1.80 per pound?
----------------------
Equations
Weight equation:::: a + b = 20 lbs
Value equation:: 1.5a + 2b = 1.8*20
===========================================
Modify for elimination:
2a + 2b = 40
1.5a+ 2b = 36
---------------------
Subtract and solve for "a":
0.5a = 4
a = 8 lbs (amt. of A candy needed)
-----
a + b = 20
b = 20-8 = 12 lbs. (amt. of B candy needed)
===================================
Cheers,
Stan H.
========================